how long will it take for $1000 to grow to $42,900 at an interest rate of 10.6% if the intrest is compounded continuously? Round the number of years to the nearest hundredth.
i got 35.46 years and 3546.11 years

which is correct? :)

Answer by stanbon(48535)

how long will it take for $1000 to grow to $42,900 at an interest rate of 10.6% if the intrest is compounded continuously? Round the number of years to the nearest hundredth.
-----------------------
A(t)=Pe^(rt)
42,900 = 1000*e^(0.106t)
Divide by 1000 to get:
e^0.106t = 42.9
Take the natural log to get:
0.106t = 3.75887..
t = 35.46 years
=========================
Cheers,
Stan H.